OpenStudy (anonymous):

Sam is throwing a baseball straight up in the air. He releases the ball 5ft. from the ground at a velocity of 95 ft./s. The acceleration due to gravity is 32 ft./s^2. This problem can be solved using the following equation. (1/2) 32t^2 - 95t-5=0 How many seconds will it take the ball to reach the ground (round to the nearest hundredth) A.5.99 Seconds B. 6.49 seconds C.7.04 seconds D.5.54 seconds

OpenStudy (anonymous):

Do you think you could help me @ash2326

OpenStudy (ash2326):

It's similar to the last question \[\frac 1 2 32t^2-95t-5=0\] or \[16t^2-95t-5=0\] \[a=16, b=-95, c=-5\] Now you need to try this @kaek98 Use quadratic formula

OpenStudy (anonymous):

Ok I just tried it on a piece of paper and I got 5.989 if that was rounded it would be 5.99

OpenStudy (ash2326):

Good :)

OpenStudy (anonymous):

can you help me on this little question? State wether the following is true: sometimes, never, or always 19+ |x| >18

OpenStudy (anonymous):

A. never B. sometimes C. always D. cannot be determined from the given information

OpenStudy (ash2326):

Subtract 19 from both the sides

OpenStudy (anonymous):

so the 19 in the beginning is gone

OpenStudy (anonymous):

where does the other 19 subtract from? the 18?

OpenStudy (ash2326):

yes you'd get \[|x|>18-19\] While you simplify this, think about absolute value function |x|, about its basic property.

OpenStudy (anonymous):

so it would be |x|> -1

OpenStudy (ash2326):

yes, now tell me is this true always, sometimes or never

OpenStudy (anonymous):

true?

OpenStudy (anonymous):

always

OpenStudy (anonymous):

true always*?

OpenStudy (ash2326):

yes, how do you say that?

OpenStudy (anonymous):

What do you mean?

OpenStudy (ash2326):

What's the reason ?

OpenStudy (anonymous):

Because x is greater than -1

OpenStudy (ash2326):

It's not given in the question!

OpenStudy (anonymous):

what?

OpenStudy (ash2326):

that x is greater than 1 hint: it's something to do with the absolute value function

OpenStudy (anonymous):

I still don't quite get it. I mean the only thing I can think of is that 19 is greater than 18 to begin with

OpenStudy (ash2326):

Absolute value is always greater than or equal to 0 \[|x| \ge 0\] for any value of x so \[|x| \ge -1\] this would be true always the left side will be a positive no.

OpenStudy (anonymous):

Ohh Okay

OpenStudy (ash2326):

Do you understand ?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Thank you

OpenStudy (ash2326):

Welcome :)